zarkam21:

Write each complex number in rectangular form. Plot and label (with a - d) each point on the polar axes below.

zarkam21:

1 attachment
zarkam21:

1 attachment
Vocaloid:

let's use a) as an example 2(cos(135) + isin(135)) you basically just need to distribute the 2 to get 2cos(135) + 2i*sin(135), then replace cos(135) and sin(135) with their actual values (use a calculator)

zarkam21:

2(cos(135) + isin(135)) = -1 + I sqrt 2

zarkam21:

= -1 + i sqrt 2 *****

zarkam21:

wait

Vocaloid:

hm 2*cos(135) = -2/sqrt(2) so -2/sqrt(2) + sqrt(2) * i should do it

Vocaloid:

(the coefficient sqrt(2) is written before the i)

zarkam21:

i AM still getting the same answer

zarkam21:

1 attachment
Vocaloid:

parentheses are important -1 + i sqrt 2 is not the same as [-1 + i]sqrt(2)

Vocaloid:

the one on top is the one written in standard form

zarkam21:

Okay so it needs to include parentheses making the answer [-1 + i]sqrt(2)

zarkam21:

solve that i get -1.41

Vocaloid:

you do not need to do anything with the coordinates once you have them -2/sqrt(2) + sqrt(2) * i is sufficient, nothing else needs to be done

Vocaloid:

similar thought process for b)

zarkam21:

3cos(120) + 3i*sin(120), 3*cos(120)=-3/sqrt3

zarkam21:

i think ;O

Vocaloid:

cos(120) = -1/2 so 3cos(120) = -3/2 don't forget about the 3*i*sin(120)

zarkam21:

-3/2 3isqrt3/2

Vocaloid:

good so -3/2 + [3*sqrt(3)/2]i is your solution

zarkam21:

cos(5pi/4) = -1/sqrt2 so 5cos(5pi/4) = -5/2 5*i*sin(5pi/4) = -5i/sqrt2

Vocaloid:

almost 5cos(5pi/4) = -5/sqrt(2) so your sol'n is -5/sqrt(2) + [-5/sqrt(2)]i

Vocaloid:

any attempts on d yet?

zarkam21:

cos(5pi/3) = 1/2 so 4cos(5pi/3) = 2 4*i*sin(5pi/3) = -2i/sqrt3

Vocaloid:

good so 2 - (-2/sqrt(3))i

Vocaloid:

wait isn't it just -2sqrt(3) i

zarkam21:

okay so c is cos(5pi/4) = -1/sqrt2 so 5cos(5pi/4) = -5/2 5*i*sin(5pi/4) = -5i/sqrt2 and d is cos(5pi/3) = 1/2 so 4cos(5pi/3) = 2 4*i*sin(5pi/3) = -2i/sqrt3i

Vocaloid:

you need to combine the values to get the complete coordinate c: -5/sqrt(2) + [-5/sqrt(2)]i d: 2 - 2sqrt(3) * i

Vocaloid:

those are your solutions